OpenStudy (anonymous):

-10=-12+4/7n 4/7 is a fraction. Can anybody help me?

OpenStudy (ranga):

-10 = -12 + 4/7 * n add 12 2 = 4/7 * n multiply by 7 14 = 4 * n divide by 4 14/4 = n n = 14/4 = 7/2 = 3 & 1/2